# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
[16, Fact 3.1] For a given test matrix $\mathsf{X}$} and outcome vector $\mathbf{y}$, let ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{S}}$ be the collection of satisfying sets, i.e. the collection of $s\in{\mathcal{S}}$ that would give rise to the test outcome $\mathbf{y}$ when using the test matrix $\mathsf{X}$}. Then, under the combinatorial prior, it holds that $(S\mid{\mathsf{X}},\mathbf{y})\sim\mathrm{Unif}({\mathcal{V}}({\mathsf{X}},\mathbf{y}))$.
Lemma 12 states that the posterior distribution of $S$ given the test outcomes and testing matrix is uniform over the sets that could have led to the observed outcome. This fact, in conjunction with the previously established high probability bounds on the number of masked items, allows us to bound $\mathbb{P}(\mathrm{suc})$ by showing that any estimate of size $(1+\eta^{+})k$ with contains the true defective set with low probability over the posterior distribution of $S$.
Let ${\mathcal{T}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{V}}({\mathsf{X}},\mathbf{y})$ be the satisfying sets that are a subset of $\hat{S}$, and further let $\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{T}}({\mathsf{X}},\mathbf{y})$ be the set of subsets $S^{\prime}$ for which all of the following conditions hold:
@@
Note that we use $\frac{1}{2}k^{1+\frac{\xi}{2}}$ in condition (iii) because it is the smaller of the two bounds we derived when we choose $\xi$ to be sufficiently small (the other bound is $\frac{n}{2k}k^{\frac{\gamma}{2}}$ with $k=o(\sqrt{n})$, and we will only need this stronger bound later when proving the second part of the theorem). Accordingly, the calculations presented in the previous part of the proof imply that the true defective set $S$ satisfies
$\mathbb{P}(S\notin\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y}))=o(1).$ (71)
Accordingly, we we can write the success probability as follows:
@@
Now suppose that for a given $({\mathsf{X}},\mathbf{y})$ the decoder’s estimate $\hat{S}=\hat{S}({\mathsf{X}},\mathbf{y})$ satisfies $\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert=N$. If $N=0$ then the fraction in (75) is trivially 0, so we henceforth assume that $N>0$. We refer to the elements of $\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})$ as being plausible for $\hat{S}$, and the elements of ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\setminus\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})$ as being implausible. The number of plausible sets is $N$, and we will show that $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by arguing that the number of implausible sets is $\omega(N)$.
Denote the plausible sets by $S_{1},\dotsc,S_{N}$ (the dependence on $({\mathsf{X}},\mathbf{y})$ is left implicit), and note that by definition $|S_{i}|=k$. We first argue that for every plausible set, there are many implausible sets obtained by swapping out a single element. Specifically, swapping out one masked defective item (in $S_{i}$) with one masked non-defective item in $n]\setminus\hat{S}$ produces an implausible set, because being plausible requires being a subset of $\hat{S}$. Note that $S_{i}\subseteq\hat{S}$, so combining $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$ gives that there are at most $\eta^{-}k$ masked non-defectives in $\hat{S}$, and hence at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ masked non-defectives in $n]\setminus\hat{S}$ (see the definition of $\widetilde{{\mathcal{T}}}$). Thus, there are at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k$ ways of performing the swap described above. For convenience, we apply $\eta^{+}=k^{o(1)}$ to simplify $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta^{+}k\geq\frac{1}{4}k^{1+\frac{\xi}{2}}$ for sufficiently large $k$, and work with this simpler lower bound.
@@
We claim that for each $S_{i}$ chosen, the number of $S_{j}$ ($j\neq i$) removed is at most $\eta^{+}k$, where we recall that $\eta^{+}=k^{o(1)}$. This is because $S_{i}$ and $S_{j}$ necessarily share k1k-1 defectives in common, and there is only one choice for the item in $S_{i}\setminus S_{j}$ (it must be $S_{i}$’s designated defective) and at most $\eta^{+}k$ choices for the item in $S_{j}\setminus S_{i}$ (it must be an element of $\hat{S}\setminus S_{i}$, where $|\hat{S}|=(1+\eta^{+})k$ and $|S_{i}|=k$).
Thus, the above procedure extracts at least $\frac{N}{\eta^{+}k+1}$ plausible sets, all of which have pairwise disjoint collections of implausible sets. As we already argued, each such collection contains at least $\frac{1}{4}k^{1+\frac{\xi}{2}}$ implausible sets, implying that $\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert\geq\frac{N}{\eta^{+}k+1}\frac{1}{4}k^{1+\frac{\xi}{2}}+N$. Since $\eta^{+}=k^{o(1)}$, this implies that $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}\leq\frac{N}{\frac{N}{\eta^{+}k+1}\frac{1}{4}k^{1+\frac{\xi}{2}}+N}=\frac{1}{\Omega\big(k^{\frac{\xi}{2}-o(1)}\big)}=o(1)$. Substituting this into (75), we obtain $\mathbb{P}(\mathrm{suc})=o(1)$ as claimed.
@@
When $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$, we have established the stronger claim that there are $\omega(\frac{n}{k})$ masked non-defective items with high probability (rather than only $\omega(k)$). We claim that this implies the same result when $\eta^{+}=\Omega(k^{\lambda})$ for $\lambda\in[0,\frac{1}{\theta}-2)$ (rather than only $\eta^{+}=k^{o(1)}$). This is because under such a choice, we have (via $k=\Theta(n^{\theta})$) that the number of possible swaps with items not in $\hat{S}$ is at least $\frac{n}{2k}k^{\frac{\gamma}{2}}-\eta^{+}k=\Theta(\frac{n}{k}k^{\frac{\gamma}{2}})$, and that $\frac{(n/k)k^{\frac{\gamma}{2}}}{\eta^{+}k}\to\infty$, which implies $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by the same reasoning as that above. ∎
@@ IV Conclusion
We have established optimal rates for group testing with one-sided approximate recovery; specifically, for SUBSET we established an algorithm whose rate matches the counting bound, and for SUPERSET we established a converse that matches the better of two existing achievability results. An immediate direction for further work would be to attain the optimal rate for SUBSET with polynomial decoding time, possibly building on the spatial coupling approach to exact and two-sided approximate recovery [16]. As discussed in Remark 1, another possible direction is to establish the best possible values of $\lambda$ that can be attained in results of the kind stated in Theorems 1 to 3.
@@
Here we provide statements of some standard concentration inequalities (e.g., see [32, Ch. 2]) that we use throughout the paper. Letting $X_{1},X_{2},\dots,X_{n}$ be a sequence of i.i.d. Bernoulli($\mu$) random variables, we have the following:
@@
    $\mathbb{P}\left(\sum_{i=1}^{n}X_{i}\geq(1+\delta)n\mu\right)\leq\exp\Big(-n\mu\big((1+\delta)\log(1+\delta)-\delta\big)\Big),$ (76)
@@
    $\mathbb{P}\left(\sum_{i=1}^{n}X_{i}\leq(1-\delta)n\mu\right)\leq\exp\Big(-n\mu\big((1-\delta)\log(1-\delta)+\delta\big)\Big).$ (77)
@@
Here we outline the proof that algorithm in the algorithm in [18] still succeeds for two-sided approximate recovery when $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ and $\lambda\in\big[0,\frac{1}{2}\big)$, as opposed to only constant $\beta\in(0,1)$. We note that a variant of this idea (with $\beta=\Theta(k^{\gamma})$ for $\gamma\in(0,1)$) has been analyzed in [29], but the details differ slightly since they consider the noisy case.
Recall the condition in (10) defining the decoder, and let $TI(\tau)$ be the mean of $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$. (Here $I(\tau)$ is a conditional mutual information, but the precise definitions of $\imath^{T}$ and $I(\tau)$ are not needed for our purposes.) The following condition on the number of tests is stated in [18, Eqn. 3.37] for parameters $\delta_{1}$ and $\{\delta_{2,\tau}\}_{\tau=\beta k+1}^{k}$:
@@
Letting $\mathcal{E}$} denote the failure event for two-sided approximate recovery, [18, Eqn. 3.38] states that if (80) holds for all $\tau=\beta k+1,\dots,k$, then the error probability is upper bounded as follows:
$\mathbb{P}({\mathcal{E}})\leq\sum_{\tau=\beta k+1}^{k}{k\choose\tau}\psi_{\tau}(T,\delta_{2,\tau})+\delta_{1},$ (81)
where $\psi_{\tau}:\mathbb{Z}\times\mathbb{R}\to\mathbb{R}$ comes from a concentration inequality given in [18, Eqn. 3.45], and is defined by
$\psi_{\tau}(T,\delta_{2,\tau})=\begin{cases}\exp\Big(-T\frac{\tau}{2k}\log 2\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)&\tau\leq\frac{k}{\log k}\\ 2\exp\left(-\frac{(\delta_{2}^{(2)}I(\tau))^{2}T}{4(8+\delta_{2}^{(2)}I(\tau))}\right)&\tau>\frac{k}{\log k}\end{cases}$ (82)
@@
We apply (81) with $\delta_{1}=\frac{1}{k}$, so that the second term is insignificant. As for each $\delta_{2,\tau}$, this is set to $\delta_{2,\tau}=\delta_{2}^{(1)}$ for $\tau\leq\frac{k}{\log k}$, and to $\delta_{2,\tau}=\delta_{2}^{(2)}$ for $\tau>\frac{k}{\log k}$, where $\delta_{2}^{(1)}$ and $\delta_{2}^{(2)}$ will be specified shortly. We can then split the summation in (81) as follows:
$\mathbb{P}({\mathcal{E}})$ $\leq\sum_{\tau=\beta k+1}^{\frac{k}{\log k}}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(1)})+\sum_{\tau=\frac{k}{\log k}+1}^{k}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(2)})+\frac{1}{k}.$ (83)
We can directly use part of the analysis in [18, Appendix B], showing that (i) (80) is satisfied for all $\tau=\beta k+1,\dots,k$, and (ii) the second sum in (83) tends to 0 for arbitrarily small $\delta_{2}^{(2)}>0$, as long as $T\geq\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Hence, to show that $\mathbb{P}({\mathcal{E}})\to 0$, it suffices to show that the first sum in (83) also decays to zero. Note that if $\beta k+1>\frac{k}{\log k}$ then the sum is empty and is trivially zero, so henceforth we assume this is not the case. We proceed to bound this first sum as follows via (82):
${\mathcal{A}}$ $\coloneqq\sum_{\tau=\beta k+1}^{\frac{k}{\log k}}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(1)})$ (84)
$=\sum_{\tau=\beta k+1}^{\frac{k}{\log k}}{k\choose\tau}\exp\Big(-T\frac{\tau}{2k}\log 2\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (85)
$\leq k\underset{\beta k+1\leq\tau\leq\frac{k}{\log k}}{\max}{k\choose\tau}\exp\Big(-T\frac{\tau}{2k}\log 2\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (86)
$=\underset{\beta k+1\leq\tau\leq\frac{k}{\log k}}{\max}\exp\Big(\log k+\log{k\choose\tau}-T\frac{\tau}{2k}\log 2\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (87)
$\leq\underset{\beta k+1\leq\tau\leq\frac{k}{\log k}}{\max}\exp\Big(\log k+\log{k\choose\tau}-T\frac{\tau}{2k}\cdot\frac{\log 2}{1+\varepsilon^{\prime}}\Big),$ (88)
where (88) follows by taking $\delta_{2}^{(1)}$ to be arbitrarily close to $1$ and introducing $\varepsilon^{\prime}>0$ that can be made arbitrarily small (as a function of $\xi$ and $\delta_{2}^{(1)}$). In order for ${\mathcal{A}}\to 0$, we require that the exponent in (88) approaches -$\infty$. For this to happen, it suffices that the following holds for all $\tau\in\big\{\beta k+1,\dotsc,\frac{k}{\log k}\big\}$:
@@
Observe that the strictest requirement on $T$ is when $\tau=\beta k$. Substituting this into (92) and recalling that $\beta=\Omega(k^{{}^{-\lambda(1-\theta)}})$ and $k=\Theta(n^{\theta})$, the bound simplifies to
$T$ $\geq 2k\big(\log_{2}k^{\lambda(\frac{1}{\theta}-1)}\big)(1+2\varepsilon^{\prime})\big(1+o(1)\big)$ (93)
@@
-C Adaptations of Existing Intermediate Results (Lemma 10 and Lemma 11)
In this section, we provide details on the modifications we make to the results in [16, 30] to obtain Lemmas 10 and 11. Throughout the section, we will use symbols with a tilde (e.g., $\tilde{n}$) to denote quantities in a “denser” scaling regime, which will then be related to “sparser” regimes with the usual notation (e.g., $n$). When the two settings both have the same value of a certain parameter, the tilde will be omitted (e.g., $k$).

### v4

@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
[10, Fact 3.1] For a given test matrix $\mathsf{X}$} and outcome vector $\mathbf{y}$, let ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{S}}$ be the collection of satisfying sets, i.e., the collection of $s\in{\mathcal{S}}$ that would give rise to the test outcome $\mathbf{y}$ when using the test matrix $\mathsf{X}$}. Then, under the combinatorial prior, the posterior distribution of $S$ satisfies $(S\mid{\mathsf{X}},\mathbf{y})\sim\mathrm{Unif}({\mathcal{V}}({\mathsf{X}},\mathbf{y}))$.
Lemma 13 states that the posterior distribution of $S$ given the test outcomes and testing matrix is uniform over the sets that could have led to the observed outcomes. This fact, in conjunction with the previously established high probability bounds on the number of masked items, allows us to bound $\mathbb{P}(\mathrm{suc})$ by showing that any estimate of size $(1+\eta)k$ with contains the true defective set with low probability over the posterior distribution of $S$.
Let ${\mathcal{T}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{V}}({\mathsf{X}},\mathbf{y})$ be the size-$k$ satisfying sets that are a subset of $\hat{S}$, and further let $\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\subseteq{\mathcal{T}}({\mathsf{X}},\mathbf{y})$ be the set of subsets $S^{\prime}$ for which all of the following conditions hold:
@@
Note that we use $\frac{1}{2}k^{1+\frac{\xi}{2}}$ in condition (iii) because it is the smaller of the two bounds we derived when we choose $\xi$ to be sufficiently small (the other bound is $\frac{n}{2k}k^{\frac{\gamma}{2}}$ with $k=o(\sqrt{n})$, and we will only need this stronger bound later when proving the second part of the theorem). Accordingly, the calculations presented in the previous part of the proof imply that the true defective set $S$ satisfies condition (iii) with probability approaching one, or equivalently
$\mathbb{P}(S\in{\mathcal{T}}({\mathsf{X}},\mathbf{Y})\setminus\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{Y}))=o(1).$ (90)
Accordingly, we can write the success probability as follows:
@@
Now suppose that for a given pair $({\mathsf{X}},\mathbf{y})$, the decoder’s estimate $\hat{S}=\hat{S}({\mathsf{X}},\mathbf{y})$ satisfies $\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert=N$. If $N=0$ then the fraction in (94) is trivially 0, so we henceforth assume that $N>0$. We refer to the elements of $\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})$ as being plausible for $\hat{S}$, and the elements of ${\mathcal{V}}({\mathsf{X}},\mathbf{y})\setminus\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})$ as being implausible. The number of plausible sets is $N$, and we will show that $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by arguing that the number of implausible sets is $\omega(N)$.
Denote the plausible sets by $S_{1},\dotsc,S_{N}$ (the dependence on $({\mathsf{X}},\mathbf{y})$ is left implicit), and note that by definition $|S_{i}|=k$. We first argue that for every plausible set, there are many implausible sets obtained by swapping out a single element. Specifically, swapping out one masked defective item (in $S_{i}$) with one masked non-defective item in $n]\setminus\hat{S}$ produces an implausible set, because being plausible requires being a subset of $\hat{S}$. Note that $S_{i}\subseteq\hat{S}$, so combining $|\hat{S}|=(1+\eta)k$ and $|S_{i}|=k$ gives that there are at most $\eta k$ masked non-defectives in $\hat{S}$, and hence at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta k$ masked non-defectives in $n]\setminus\hat{S}$ (see the definition of $\widetilde{{\mathcal{T}}}$). Thus, there are at least $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta k$ ways of performing the swap described above. For convenience, we apply $\eta=k^{o(1)}$ to simplify $\frac{1}{2}k^{1+\frac{\xi}{2}}-\eta k\geq\frac{1}{4}k^{1+\frac{\xi}{2}}$ for sufficiently large $k$, and work with this simpler lower bound.
@@
We claim that for each $S_{i}$ chosen, the number of $S_{j}$ ($j\neq i$) removed is at most $\eta k$, where we recall that $\eta=k^{o(1)}$. This is because $S_{i}$ and $S_{j}$ necessarily share k1k-1 items in common, and there is only one choice for the item in $S_{i}\setminus S_{j}$ (it must be $S_{i}$’s designated defective) and at most $\eta k$ choices for the item in $S_{j}\setminus S_{i}$ (it must be an element of $\hat{S}\setminus S_{i}$, where $|\hat{S}|=(1+\eta)k$ and $|S_{i}|=k$).
Thus, the above procedure extracts at least $\frac{N}{\eta k+1}$ plausible sets, all of which have pairwise disjoint collections of implausible sets. As we already argued, each such collection contains at least $\frac{1}{4}k^{1+\frac{\xi}{2}}$ implausible sets, implying that $\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert\geq\frac{N}{\eta k+1}\frac{1}{4}k^{1+\frac{\xi}{2}}+N$. Since $\eta=k^{o(1)}$, this implies that $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}\leq\frac{N}{\frac{N}{\eta k+1}\frac{1}{4}k^{1+\frac{\xi}{2}}+N}=\frac{1}{\Omega\big(k^{\frac{\xi}{2}-o(1)}\big)}=o(1)$. Substituting this into (94), we obtain $\mathbb{P}(\mathrm{suc})=o(1)$ as claimed.
@@
When $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$, we have established the stronger claim that there are $\omega(\frac{n}{k})$ masked non-defective items with high probability (rather than only $\omega(k)$). We claim that this implies the same result when $\eta=\Omega(k^{\lambda})$ for $\lambda\in[0,\frac{1}{\theta}-2)$ (rather than only $\eta=k^{o(1)}$). This is because under such a choice, we have (via $k=\Theta(n^{\theta})$) that the number of possible swaps with items not in $\hat{S}$ is at least $\frac{n}{2k}k^{\frac{\gamma}{2}}-\eta k=\Theta(\frac{n}{k}k^{\frac{\gamma}{2}})$, and that $\frac{(n/k)k^{\frac{\gamma}{2}}}{\eta k}\to\infty$, which implies $\frac{\lvert\widetilde{{\mathcal{T}}}({\mathsf{X}},\mathbf{y})\rvert}{\lvert{\mathcal{V}}({\mathsf{X}},\mathbf{y})\rvert}=o(1)$ by the same reasoning as that above. ∎
@@ IV Conclusion
We have established optimal rates for group testing with one-sided approximate recovery; specifically, for SUBSET we established an algorithm whose rate matches the counting bound, and for SUPERSET we established a converse that matches the better of two existing achievability results.
An immediate direction for further work would be to attain the optimal rate for SUBSET with polynomial decoding time, possibly building on the spatial coupling approach to exact and two-sided approximate recovery [10]. While such an approach gives us the required starting point of a two-sided approximate recovery estimate, the subsequent analysis in Section III-A may be difficult to adapt to the more complicated test design. Perhaps more fundamentally, even given the initial estimate, our subsequent search step is still over $n^{\omega(1)}$ candidate sets, so the question remains of how to maintain its performance with polynomial runtime.
As discussed in Remark 1, another possible direction is to establish the best possible values of $\lambda$ that can be attained in results of the kind stated in Theorems 1 to 3.
@@
Here we provide statements of some standard concentration inequalities (e.g., see [6, Ch. 2]) that we use throughout the paper. Letting $X_{1},X_{2},\dots,X_{n}$ be a sequence of i.i.d. Bernoulli($\mu$) random variables, we have the following:
@@
    $\mathbb{P}\left(\sum_{i=1}^{n}X_{i}\geq(1+\delta)n\mu\right)$
    $\qquad\qquad\leq\exp\Big(-n\mu\big((1+\delta)\log(1+\delta)-\delta\big)\Big),$ (95)
@@
    $\mathbb{P}\left(\sum_{i=1}^{n}X_{i}\leq(1-\delta)n\mu\right)$
    $\qquad\qquad\leq\exp\Big(-n\mu\big((1-\delta)\log(1-\delta)+\delta\big)\Big).$ (96)
@@
Here we outline the proof that algorithm in the algorithm in [28] still succeeds for two-sided approximate recovery when $\beta=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ and $\lambda\in\big[0,\frac{1}{2}\big)$, as opposed to only constant $\beta\in(0,1)$. We note that a variant of this idea (with $\beta=\Theta(k^{\gamma-1})$ for $\gamma\in(0,1)$) has been analyzed in [30], but the details differ slightly since they consider the noisy case.
Recall the condition in (10) defining the decoder, and let $TI(\tau)$ be the mean of $\imath^{T}\left({\mathsf{X}}_{s_{\mathrm{dif}}};\mathbf{y}\mid{\mathsf{X}}_{s_{\mathrm{eq}}}\right)$. (Here $I(\tau)$ is a conditional mutual information, but the precise definitions of $\imath^{T}$ and $I(\tau)$ are not needed for our purposes.) The following condition on the number of tests is stated in [28, Eqn. 3.37] for parameters $\delta_{1}$ and $\{\delta_{2,\tau}\}_{\tau=\lfloor\beta k+1\rfloor}^{k}$:
@@
Letting $\mathcal{E}$} denote the failure event for two-sided approximate recovery, [28, Eqn. 3.38] states that if (99) holds for all $\tau=\lfloor\beta k+1\rfloor,\dots,k$, then the error probability is upper bounded as follows:
$\mathbb{P}({\mathcal{E}})\leq\sum_{\tau=\lfloor\beta k+1\rfloor}^{k}{k\choose\tau}\psi_{\tau}(T,\delta_{2,\tau})+\delta_{1},$ (100)
where $\psi_{\tau}:\mathbb{Z}\times\mathbb{R}\to\mathbb{R}$ comes from a concentration inequality given in [28, Eqn. 3.45], and is defined by
$\psi_{\tau}(T,\delta_{2,\tau})=$
$\begin{cases}\exp\Big(-T\frac{\tau}{2k}\log 2\Big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})\\ \qquad\qquad\qquad\qquad\quad+\delta_{2}^{(1)}\Big)(1-\xi)\Big)&\hskip-8.1pt\tau\leq\lfloor\frac{k}{\log k}\rfloor\\ 2\exp\left(-\frac{(\delta_{2}^{(2)}I(\tau))^{2}T}{4(8+\delta_{2}^{(2)}I(\tau))}\right)&\hskip-8.1pt\tau>\lfloor\frac{k}{\log k}\rfloor\end{cases}$ (101)
@@
We apply (100) with $\delta_{1}=\frac{1}{k}$, so that the second term is insignificant. As for each $\delta_{2,\tau}$, this is set to $\delta_{2,\tau}=\delta_{2}^{(1)}$ for $\tau\leq\lfloor\frac{k}{\log k}\rfloor$, and to $\delta_{2,\tau}=\delta_{2}^{(2)}$ for $\tau>\lfloor\frac{k}{\log k}\rfloor$, where $\delta_{2}^{(1)}$ and $\delta_{2}^{(2)}$ will be specified shortly. We can then split the summation in (100) as follows:
$\mathbb{P}({\mathcal{E}})$ $\leq\sum_{\tau=\lfloor\beta k+1\rfloor}^{\lfloor\frac{k}{\log k}\rfloor}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(1)})$
$+\sum_{\tau=\lfloor\frac{k}{\log k}\rfloor+1}^{k}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(2)})+\frac{1}{k}.$ (102)
We can directly use part of the analysis in [28, Appendix B], showing that if $T\geq\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$, then (i) (99) is satisfied for all $\tau=\lfloor\beta k+1\rfloor,\dots,k$, and (ii) the second sum in (102) tends to 0 for arbitrarily small $\delta_{2}^{(2)}>0$. Hence, to show that $\mathbb{P}({\mathcal{E}})\to 0$, it suffices to show that the first sum in (102) also decays to zero. Note that if $\beta k+1>\frac{k}{\log k}$ then the sum is empty and is trivially zero, so henceforth we assume this is not the case. We proceed to bound this first sum as follows via (101):
${\mathcal{A}}$ $\coloneqq\sum_{\tau=\lfloor\beta k+1\rfloor}^{\lfloor\frac{k}{\log k}\rfloor}{k\choose\tau}\psi_{\tau}(T,\delta_{2}^{(1)})$ (103)
$=\sum_{\tau=\lfloor\beta k+1\rfloor}^{\lfloor\frac{k}{\log k}\rfloor}{k\choose\tau}\exp\Big(-T\frac{\tau}{2k}\log 2$
$\qquad\qquad\times\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (104)
$\leq k\underset{\lfloor\beta k+1\rfloor\leq\tau\leq\lfloor\frac{k}{\log k}\rfloor}{\max}{k\choose\tau}\exp\Big(-T\frac{\tau}{2k}\log 2$
$\qquad\qquad\times\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (105)
$=\underset{\lfloor\beta k+1\rfloor\leq\tau\leq\lfloor\frac{k}{\log k}\rfloor}{\max}\exp\Big(\log k+\log{k\choose\tau}-T\frac{\tau}{2k}$
$\qquad\times\log 2\big((1-\delta_{2}^{(1)})\log(1-\delta_{2}^{(1)})+\delta_{2}^{(1)}\big)(1-\xi)\Big)$ (106)
$\leq\underset{\lfloor\beta k+1\rfloor\leq\tau\leq\lfloor\frac{k}{\log k}\rfloor}{\max}\exp\Big(\log k+\log{k\choose\tau}$
$\qquad\qquad\qquad\qquad\qquad\qquad-T\frac{\tau}{2k}\cdot\frac{\log 2}{1+\varepsilon^{\prime}}\Big),$ (107)
where (107) follows by taking $\delta_{2}^{(1)}$ to be arbitrarily close to $1$ and introducing $\varepsilon^{\prime}>0$ that can be made arbitrarily small (as a function of $\xi$ and $\delta_{2}^{(1)}$). In order for ${\mathcal{A}}\to 0$, we require that the exponent in (107) approaches -$\infty$. For this to happen, it suffices that the following holds for all $\tau\in\big\{\lfloor\beta k+1\rfloor,\dotsc,\lfloor\frac{k}{\log k}\rfloor\big\}$:
@@
Observe that the strictest requirement on $T$ is when $\tau=\lfloor\beta k+1\rfloor$. Substituting this into (111) and recalling that $\beta=\Omega((\frac{k}{n})^{\lambda})$, the bound simplifies to
$T$ $\geq 2k\Big(\log_{2}\Big(\Big(\frac{n}{k}\Big)^{\lambda}\Big)\Big)(1+2\varepsilon^{\prime})\big(1+o(1)\big)$ (112)
@@
-C Adaptations of Existing Intermediate Results (Lemmas 10-12)
In this section, we provide details on the modifications we make to the results in [10, 5] to obtain Lemmas 10-12. Throughout the section, we will use symbols with a tilde (e.g., $\tilde{n}$) to denote quantities in a “denser” scaling regime, which will then be related to “sparser” regimes with the usual notation (e.g., $n$). When the two settings both have the same value of a certain parameter, the tilde will be omitted (e.g., $k$).